Compute the population mean 2,5,7,9,10,11,12


1
Expert's answer
2021-05-21T03:49:00-0400

Solution:

Mean"=\dfrac{\Sigma f_i}{n}=\dfrac{2+5+7+9+10+11+12}7"

"=\dfrac{56}7=8"